OpenStudy (anonymous):

What is the solution to the system? Line 1 x y –1 –2 1 2 Line 2 x y –2 10 4 1 (_____,_____) i think it is 2,4

OpenStudy (anonymous):

@livias.random is this right?

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

i am just gonna check again real quick one sec

OpenStudy (anonymous):

yep by my equation i am right what did u get @livias.random ?

OpenStudy (anonymous):

m = (-2-2)/(-1-1) m = -4/-2 m = 2 (y-2) = 2(x-1) y-2 = 2x-2 y = 2x - m = (10-1)/(-2-4) m = 9/-6 m = -3/2 (y-1) = -3/2(x-4) y-1 = -3/2x+6 y = -3/2x+7 - 2x = -3/2x+7 2x+3/2x = 7 7/2x = 7 x = 2 - y = 2x y = 2(2) y = 4 - (2,4) So yep! I think you're right :)

OpenStudy (anonymous):

yay! here is medal and my testimony i was gonna give u earlier!

OpenStudy (anonymous):

i hope u like!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Aw so sweet :)